OpenStudy (anonymous):

Jordan poured 750 milliliter of water into a vase. What fractional part of a liter is 750 milliliters?

OpenStudy (anonymous):

750 milliliters = 0.75 liters Do you know how to turn that into a fraction?

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

L = mL/1000.0 L = Liter mL = Milliliter

OpenStudy (anonymous):

So, insert what you know: L = 750/1000.0 L = 0.75 0.75 = 75/100 = 3/4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No problem, this is my job! :-)